OpenStudy (kpop4life123):

If y = 7x – 5, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs?

OpenStudy (kpop4life123):

{(–5, 0), (9, 2), (–26, –3)} {(2, 7), (1, 6), (3, 13)} {(0, –5), (2, 9), (–3, –26)} {(1, 3), (6, 18), (8, 15)}

OpenStudy (♪chibiterasu):

Plug in x for the 'x' in the equation and y for the 'y' in the equation. Solve, if they are equal, it is true.

OpenStudy (♪chibiterasu):

You don't have to do all of them, just try the first one of each.

OpenStudy (kpop4life123):

i dont understand on how to do it

OpenStudy (♪chibiterasu):

Let's try the first one: y = 7x – 5 {(–5, 0), (9, 2), (–26, –3)} Let's take (-5, 0) and plug in. x = -5 y = 0 0 = 7(-5) – 5 0 = -35 - 5 0 = -40 Since they are not equal, it is false, making A wrong.

OpenStudy (kpop4life123):

ohhhh ok

OpenStudy (♪chibiterasu):

Let me know what you get :)

OpenStudy (kpop4life123):

u can go ill try doing all of them and ill tag u wen i find the answer :)

OpenStudy (♪chibiterasu):

ㅇㅋ :)